### v1

# From Conversation to Automation: Leveraging Large Language Models to Analyze Strategies in Problem Solving Therapy

Abstract
Problem-solving therapy (PST) is a structured psychological approach that helps individuals manage stress and resolve personal issues by guiding them through problem identification, solution brainstorming, decision-making, and outcome evaluation. As mental health care increasingly integrates technologies like chatbots and large language models (LLMs), understanding how PST can be effectively automated is important. This study leverages anonymized therapy transcripts to analyze and classify therapeutic interventions using various LLMs and transformer-based models. Our results show that GPT-4o achieved the highest accuracy (0.76) in identifying PST strategies, outperforming other models. Additionally, we introduced a new dimension of communication strategies that enhances the current PST framework, offering deeper insights into therapist-client interactions. This research demonstrates the potential of LLMs to automate complex therapeutic dialogue analysis, providing a scalable, efficient tool for mental health interventions. Our annotation framework can enhance the accessibility, effectiveness, and personalization of PST, supporting therapists in real-time with more precise, targeted interventions.
From Conversation to Automation: Leveraging Large Language Models to Analyze Strategies in Problem Solving Therapy
Elham Aghakhani1, Lu Wang2, Karla T. Washington3 George Demiris4, Jina Huh-Yoo2, Rezvaneh Rezapour1 1Drexel University, {ea664,sr3563}@drexel.edu 2Stevens Institute of Technology, {lwang97,jhuhyoo}@stevens.edu 3Washington University, kwashington@wustl.edu 4University of Pennsylvania, gdemiris@nursing.upenn.edu
1 Introduction
In recent years, the increasing complexities of modern life have significantly influenced the prevalence of developing mental health conditions. According to the American Psychological Association (APA), the COVID-19 pandemic, global conflicts, racism, inflation, and climate disasters have led to widespread collective trauma among Americans, with increased chronic illnesses in those aged 35 to 44 and a significant rise in mental health diagnoses (APA, 2023). In response, the World Health Organization (WHO) underscores the urgent transformation of global mental health care (WHO, 2022). Building on this, Problem-Solving Therapy (PST) has emerged as a pivotal therapeutic approach to help individuals manage these stressors and challenges more effectively (Song and Sun, 2019). Rooted in the principle that improving problem-solving skills reduces psychological distress, PST fosters resilience and well-being (Arean et al., 2008). By guiding individuals to address immediate problems with practical strategies (Figure 1), PST strengthens coping mechanisms and mitigates negative impacts on both mental and physical health (Gellis and Kenaley, 2008).
Figure 1: A real-life example of therapist and client exchanges in a problem solving therapy session.
As rapidly evolving trends continue to shape the critical high-stakes area of mental health, there is currently a lack of systematic methods for exploring and understanding PST between clients and therapists Chiu et al. (2024). This gap becomes particularly relevant with the increasing prevalence of mental health chatbots, especially those deployed using large language models (LLMs) Lawrence et al. (2024); Cho et al. (2023). Before these models are widely adopted, it is crucial to examine how mental health as well as PST strategies are implemented in practice. Previous research highlighted the variability in therapists’ application of PST techniques, which are often tailored to individual client needs and contexts Andrade-Arenas et al. (2024). In PST sessions, therapists follow a structured process: identifying the problem, brainstorming solutions, making decisions, and evaluating outcomes—all within a supportive framework. This approach equips clients with PS skills to tackle future challenges, enhancing their sense of control and overall coping abilities (Malouff et al., 2007). Therefore, while the new technological solutions hold promises, it is essential to understand how PST can be effectively automated, ensuring these tools are responsive and sensitive to the varied situations of clients Izadi and Forouzanfar (2024).
Previous work leveraged LLMs for analyzing therapist-client interactions Lin et al. (2024); Demszky et al. (2023). For instance, Lee et al. used LLMs, particularly ChatGPT, to identify utterance-level features in therapist-client conversations to classify counselor’s utterances. In PST, such research in automated annotations is still lacking due to limited data and costly domain expert annotation. Automating the annotation of various cognitive therapy techniques is crucial for developing efficient models, like LLMs, that can navigate the complexity of therapeutic dialogues and identify effective interventions. This capability could enable real-time suggestions to therapists or direct interventions in automated systems.
Our work addresses this gap by using LLMs to automate the annotation of PST, leveraging both closed and open-weight models to categorize therapist utterances. This approach allows us to efficiently analyze therapist-client interactions and enhances the scalability and precision of mental health interventions through automated systems. More specifically, we used anonymized real-life 240 PST conversations between family caregivers and their therapists, consisting of 68,306 dialogue exchanges. We applied both open- and closed-weight LLMs to automatically annotate therapist utterances with strategies based on existing and adapted codebooks of PST and communication strategies. Our results show that using LLMs for analyzing PST-related texts is efficient, as they achieve high accuracy for annotating PST strategies in therapeutic conversations. Additionally, our analysis is validated by examining the specific strategies therapists employed across multiple client sessions, highlighting how these strategies evolved as therapy progressed. Our approach not only facilitates a deeper understanding of therapist-client interactions but also supports the development of more responsive and adaptive mental health interventions by enabling the real-time identification and application of effective therapeutic techniques.
Our work contributes to the field of NLP and mental health care by introducing a framework for automating complex therapeutic dialogue analysis, providing a scalable, efficient tool for mental health interventions. In addition, using anonymized transcripts from real-world therapy sessions, our results advance PST research and demonstrate the transformative potential of LLMs in mental health care. Finally, adding a new dimension of communication strategies enables a deeper understanding of therapist-client interactions, thereby refining the existing PST framework and enhancing the potential for more targeted therapeutic interventions.
2 Related Work
2.1 LLMs for Mental Health Care
LLMs have been applied for mental health care, leveraging distinct features such as conversational capabilities Ma et al. (2023); Zheng et al. (2023), content generation Yang et al. (2024); Smith et al. (2023), and detection of information patterns Posner et al. (2011); Mauriello et al. (2021). Based on 34 papers out of 313 publications from six databases Hua et al. (2024) identified three key application areas of LLMs in mental health care: 1) conversational agents for digital interventions, 2) resource enrichment for training and technical development, and 3) diagnosis of mental disorders and their sources. LLMs have been integrated into conversational agents (Bendig et al., 2022) and used for personal digital companions (Ma et al., 2023), emotional support (Zheng et al., 2023), and online counseling (Sharma et al., 2023; Filienko et al., 2024).Researchers applied LLMs to generate explanations for mental health analysis Yang et al. (2024), help people learn mindfulness Kumar et al. (2023), and create hypothetical case vignettes to social psychiatry (Smith et al., 2023). LLMs have also been used in classification of subtypes of suicide (Posner et al., 2011), and identification of sources of stress span (Mauriello et al., 2021).
However, gaps remain in previous research in prioritizing privacy, safety, and ethics (Hua et al., 2024; Wang et al., 2021a; Goel et al., 2023). For example, among 36 datasets identified in Hua et al. (2024), 55.6% were collected from social media, 13.9% were simulated by clinicians, crowd workers, and LLMs, and only 8.3% came from real-life counseling conversations. Moreover, labels in existing datasets are typically not validated by experts (Hua et al., 2024; Harrigian et al., 2020). The absence of relevant evaluation frameworks brings further challenges. Every cognitive and interpersonal therapy technique has unique, specialized structure and requirement for evaluating its fidelity of delivery(Sharma et al., 2023; Mehta et al., 2022). Examining how mental health professionals deliver interventions through therapy techniques in different contexts is essential for addressing these gaps and improving professionals’ practices.
2.2 LLMs as (Co-)Annotators
LLMs have been increasingly used for annotation since acquiring high-quality labeled data is time-consuming and costly when relying solely on human experts for annotation Wang et al. (2021b, 2024); Bouzoubaa et al. (2024b). Methods such as zero-shot, without any in-domain training Liu et al. (2024); Reynolds and McDonell (2021); Bouzoubaa et al. (2024a) and few-shot learning, incorporating examples of the task Brown et al. (2020) have been used to optimize prompts Reynolds and McDonell (2021); Goel et al. (2023). Additional research showed LLMs outperformed crowd workers on annotation tasks including relevance, stance, topics, frame detection, and categorizing sentence segments Gilardi et al. (2023); He et al. (2023); Törnberg (2023). Wang et al. (2021b) found that GPT-3 cost 50% to 96% less than human for annotation. On the other hand, Zhu et al. (2023) and Ziems et al. (2024) showed that the performance of LLMs can vary substantially depending on tasks, datasets, and labels, claiming the irreplaceableness of human annotators. Wang et al. (2021b) and Li (2024) found that human-LLM collaborative approaches—either through human annotation of LLM-labeled data or combining both annotations—outperformed LLM-only or human-only approaches. Li et al. (2023c) further explored the allocation between humans and LLMs for annotation and contributed to a paradigm using uncertainty to estimate LLMs’ annotation capability. Researchers also developed human-LLM collaborative annotation systems to support the annotations Kim et al. (2024); Tang et al. (2024).
Researchers in healthcare explored a human-LLM collaboration using prompt engineering to extract medication information and found that LLMs accelerated this process with accuracy comparable to trained medical NLP annotators Goel et al. (2023). They also demonstrated the potential of a fine-tuned T5-based model for scaling data collection by generating annotation scores from interview transcripts for Bipolar Disorder and Schizophrenia Aich et al. (2024).
The studies so far, however, show limited progress on how LLMs perform as annotators in PST-related therapy. Investigating their performance on real-world, closed-domain data, like medical records, could offer valuable insights into their annotation capabilities. This experimentation may also highlight their potential beyond conversational agents, improving clinical workflows in mental health therapy delivery.
3 Method
3.1 Data
Data Collection. [Anonymized collaborator] provided transcriptions of 240 PST sessions between 152 family caregivers of older adults in senior living facilities and their therapists. Each caregiver attended up to three therapy sessions, but not every session was recorded and transcribed due to technical and legal constraints. We first ensured the anonymity of these transcripts by removing any remaining information that could potentially reveal the identity of the clients or therapists. We then converted the transcripts into 68,306 dialogue exchanges between clients and therapists. An example of two exchanges is shown in Figure 1. We then separated the utterances [^1] An utterance is a continuous piece of speech, by one person, before or after which there is silence on the part of the person: https://en.wikipedia.org/wiki/Utterance made by therapists, which amounted to 34,273 utterances. After consulting with the domain experts and to ensure high quality and relevance, we excluded utterances shorter than 5 words, such as “hmmm”, “okay”, “well”, and “oh good” assuming that longer ones would provide more meaningful content for our study. This process resulted in 14,417 utterances for further analysis. The average word count for these utterances was 29.78 with a standard deviation of 42.72.
Problem-Solving Therapy Codebook. The annotation codebook was developed based on the core strategies derived from the ADAPT model of PST, which includes five steps: 1) developing a positive Attitude, 2) Defining the problem and setting goals, 3) being creative and finding Alternative solutions, 4) Predicting how the solutions will work and making a plan, and 5) Trying out the solution plan Demiris et al. (2019). We refer to these as the ‘Problem-Solving (PS) core’ category. Upon further exploration of our data and various frameworks and codebooks for therapy conversations and strategies Li et al. (2023a); Ribeiro et al. (2013); Wang et al. (2023); Ghamari Kivi and Jamshiddoust Miyanroudi (2022), our domain experts suggested adding a new category of four strategies to this list, called ‘Communication’ strategies, to enhance the comprehensiveness of the analysis. These strategies are further explained in Table 6.
Data Annotation. To develop an evaluation set, we randomly selected 500 therapist utterances for detailed annotation. One domain expert alongside one researcher from our team performed the annotation. The annotators were instructed to select at most one strategy from the ‘PS core’ category and one strategy from the ‘Communication’ category, if applicable. If none of the strategies were suitable, we labeled it as ‘None’. We used Cohen’s Kappa to assess annotators’ agreement across different classes of PST strategies. Our findings show that the agreement scores ranged from 0.69 to 0.88, indicating a substantial agreement among the annotators, reflecting a reasonable level of consistency in identifying the strategies. The highest agreement was observed in classes that were more straightforward to identify, such as ‘Generating Alternative Solutions’, whereas the lowest agreement occurred in classes that were potentially more subjective or nuanced in nature, e.g., ‘Therapeutic Engagement’.
3.2 Classification of PST Strategies
LLM Prompting. LLMs, such as GPT-4, have been used for data annotation, offering capabilities for understanding and categorizing diverse datasets (Li et al., 2023a; Kuzman et al., 2023; Hoes et al., 2023). In this work, we employed these models to annotate therapist utterances, aiming to identify the PST strategy used in each case. Specifically, we used two versions of OpenAI’s models: gpt-4o-2024-08-06 (OpenAI, ) (referred to as GPT-4o) and gpt-4-0125-preview (OpenAI, 2024) (referred to as GPT-4), as well as two open-source models, Llama-3.1-70B-Instruct (Meta, ) (referred to as Llama) and Yi-1.5-34B (Young et al., 2024) (referred to as Yi). All models were deployed with a zero-temperature setting to ensure deterministic output, enhancing the consistency of our annotations (See Appendix C for hyperparameters). Using this approach, we compared the efficiency and accuracy of these models against our 500 annotated utterances. We used two approaches to test how well LLMs can annotate data:
    No Context: We provided specific instructions to LLMs, followed by definitions of each strategy and a few examples (one for each class). We then presented each therapist’s utterance for classification, without providing any additional conversational context.
    Two Previous Utterances as Context: In addition to instructions, strategy definitions, and examples, we provided LLMs with two previous statements from the conversation—one from the client and one from the therapist. Using this context, we evaluated how LLMs perform when conversation history is added.
The labeling process with LLMs followed the same approach as with the annotators. In each prompt, we asked the models to select one strategy from the ‘PS core’ and ‘Communication’ categories or return ‘None’ if none applied. Prompts are available in Appendix D. We ran each LLM five times to test their stability for annotation. To evaluate the consistency of the annotation, we calculated the entropy of the labels generated by the LLMs (Li et al., 2023b). Entropy (Eq. (1)) measures the degree of uncertainty or disorder within a dataset. The higher the entropy value, the more uncertain the LLMs’ responses are.
$u_{i}=-\sum_{j=1}^{k}P_{\theta}(a_{ij}\mid p_{ij})\ln P_{\theta}(a_{ij}\mid p_{ij})$ (1)
where $P_{\theta}(a_{ij}\mid p_{ij})$ is calculated as the frequency of a certain prediction among all predictions.
Model Training via LLM Labels. In addition to using LLM-based models, we also tested the performance of three transformer-based models to classify the utterances in our data. Using a subset of annotated data with the best performing LLM, we fine-tuned DeBERTa He et al. (2020), MentalBERT Ji et al. (2022), and FlanT5 Chung et al. (2024). Model details are provided in Appendix C. The choice of these models was driven by the need to explore a diverse range of transformer architectures while also including models tailored to mental health-related tasks. We fine-tuned each model for 3 epochs with a learning rate of 2e-5. Given that our annotation framework was based on two dimensions, ‘PS core’ and ‘Communication’ strategies, we fine-tuned two separate models for each dimension. Finally, we tested the performance of each fine-tuned model using our evaluation set of 500 human-annotated utterances.
4 Result
4.1 Performance of LLMs in Annotation
Table 1 shows the weighted F1 scores of four LLMs, including two proprietary models (GPT-4 and GPT-4o) and two open-source models (Llama-3.1 and Yi-1.5). We evaluate the models’ performance using our evaluation set (500 annotated utterances) for both ‘No context’ and ‘With context’ conditions. GPT-4o achieved the highest F1 score (0.76) with ‘No context’, outperforming all other models. GPT-4 followed closely with F1 score of 0.68. All models showed lower F1 scores when provided with additional context (i.e., the two previous utterances in the conversation). For GPT-4, the performance decreased from 0.68 to 0.64 when context was included. GPT-4o experienced a more significant drop, from 0.76 to 0.61. Similarly, for the smaller open-source models, Llama-3.1 and Yi-1.5, performance dropped to F1 scores of 0.50 (See Appendix E.1 for more details).
Table 2 summarizes the entropy results across five runs for each LLM. Since entropy values were the same for both ‘No context’ and ‘With context,’ we report only one set of results. The entropy values across all models were low enough to ensure consistency.Both GPT-4 and GPT-4o had higher mean entropy than Llama-3.1 and Yi-1.5, with GPT-4o slightly more consistent (0.035 vs. 0.042). Although GPT models showed higher variability (standard deviations of 0.150 and 0.130), their higher F1 scores indicate this uncertainty had no impact on the overall performance of the models. In contrast, Llama-3.1 and Yi-1.5 had much lower entropy, with Yi-1.5 showing complete consistency (0.00), but this may reflect overconfidence and poor generalization, as suggested by the lower F1 score.
Overall, the classification result shows that GPT-4o performed best without any context. Precision, recall, and F1 scores for each PST strategy are shown in Table 3. The model performed well in classifying ‘PS core’ strategies, with balanced scores across strategies like ‘Problem-Solving Positive Mindset’ and ‘Defining Problems and Goals’. For communication, it excelled in identifying ‘Therapeutic Engagement’ and ‘Problem-Solving Test Review’, while ‘Social Courtesies’ and ‘Session Management’ showed slightly lower precision. As a result of these experiments, we selected GPT-4o to annotate the entire dataset of therapist utterances. Given its low mean entropy (0.035), we only ran the model once for full annotation.
Model No context With context
GPT4 0.68 0.64
GPT4o 0.76 0.61
Llama-3.1 0.61 0.50
Yi-1.5 0.56 0.50
Table 1: F1 Score (Weighted Average) Comparison Between LLMs
Model Mean Entropy Std Entropy
GPT-4 0.042 0.150
GPT-4O 0.035 0.130
Llama 0.016 0.091
Yi-1.5 0.000 0.000
Table 2: Comparison of Mean and Standard Deviation of Entropy across LLMs
Strategies Precision Recall f1-score #Support
PS Core Problem-solving Positive Mindset 0.85 0.85 0.85 39
Defining Problems and Goals 0.86 0.73 0.79 66
Generating Alternative Solutions 0.83 0.76 0.79 36
Outcome Prediction and Planning 0.86 0.82 0.84 33
Trying Out Solution Plan 0.67 1 0.80 18
Communication Social Courtesies 0.59 0.94 0.73 25
Session Management 0.58 0.81 0.68 64
Therapeutic Engagement 0.69 0.95 0.80 130
Problem-Solving Test Review 1 0.86 0.92 24
None 0.88 0.48 0.62 90
weighted average 0.77 0.79 0.76 525
Table 3: Precision, Recall, and F1-Score for Communication and PS Strategies in GPT4o (without context)
4.2 Model Fine-Tuning Results
After annotating the entire dataset of 14,417 therapist utterances with GPT-4o, we randomly sampled 5,000 utterances and fine-tuned three transformed-based models. We used the 500 utterances annotated by human experts as the evaluation set. Table 4 shows the average weighted F1 scores for DeBERTa, MentalBERT, and FLAN-T5 on ‘PS core’ and ‘Communication’ strategies. MentalBERT achieved the highest F1 score (0.78) for ‘PS core’ strategies, while DeBERTa performed best on ‘Communication’ (0.73). Both models tied for overall performance with an F1 score of 0.68, while FLAN-T5 slightly trailed at 0.67. These results highlight DeBERTa’s strength in ‘Communication’ annotation and MentalBERT’s advantage in ‘PS core’ (See Appendix E.2).
Model PS Commu. Combined
DeBERTa 0.77 0.73 0.68
MentalBERT 0.78 0.69 0.68
FLAN-T5 0.68 0.69 0.67
Table 4: Average weighted F1 scores for PS Core (PS), Communication (Commu.), and overall across models.
Strategies Top Bigrams Top LIWC Distribution
Defining Problems and Goals (goals, around), (defining, problem), (facts, around), (obstacles, goals), (clear, language), (meeting, goals), (solving, important), (one, goals) reward 33.48%(1,743)
Generating Alternative Solutions (brainstorming, ideas), (strategies, tactics), (defer, judgment), (leads, quality), (quantity, leads), (three, things), (variety, ideas), (lots, ideas), (good, thing), (big, idea) insight, curiosity 20.03%(1,043)
PS Core Problem-solving Positive Mindset (strong, emotion), (way, feel), (affects, way), (healthy, thinking), (make, mistakes), (positive, attitude), (things, change), (thinking, rules), (really, helpful), (use, feelings), (try, balance), (problems, challenges), (feelings, adaptively) emotion, feeling 11.75%(612)
Outcome Prediction and Planning (social, consequences), (personal, consequences), (positive, negative), (solve, problem), (pros, cons), (last, time), (personal, social), (rough, screening), (would, affect) risk, focusfuture 10.18%(530)
Trying Out Solution Plan (positive, negative), (see, works), (make, sense), (really, hard), (go, back), (solve, problem), (really, important), (effort, put), (get, support), (extra, support), (talk, trying) focuspast, focusfuture 2.97%(155)
Therapeutic Engagement (really, good), (makes, sense), (right, right), (last, time), (oh, gosh), (good, good), (make, sure), (might, look), (one, things), (goals, around) focuspresent 52.19%(5,918)
Communication Session Management (next, time), (time, meet), (last, time), (defining, problem), (alternative, solutions), (next, week), (let, see), (going, talk), (time, talk), (brainstorming, ideas) focusfuture 15.80%(1,792)
Social Courtesies (thank, much), (look, forward), (thank, time), (talk, next), (taking, time), (next, time), (good, good), (thank, appreciate), (take, care), (really, helpful) emotion, prosocial 8.69%(986)
Problem-Solving Test Review (problem, solving), (positive, rational), (impulsive, avoidant), (positive, negative), (high, positive), (rational, impulsive), (negative, orientation), (negative, rational), (approach, problem), (one, questionnaires), (positive, orientation), (five, dimensions) insight 1.72%(196)
None - - 21.57%(3,111)
Table 5: Overview of Strategies and Their Most Frequent Bigrams and LIWC features with Distribution
4.3 Characteristics of the PST Data
Language Patterns in Therapeutic Strategies. Table 5 shows the distribution of therapeutic strategies and top bigrams in our labeled dataset by GPT4o. To gain deeper insights into the linguistic features of each strategy, we further analyzed the labeled utterances using the Linguistic Inquiry and Word Count (LIWC) Boyd et al. (2022). Among the ‘PS core’ strategies, ‘Defining Problems and Goals’ is the most prevalent (33.48%). This highlights the therapists’ emphasis on helping clients clarify their objectives, with words like “defining problem” and “solving important”. This strategy is further supported by LIWC features such as ‘reward,’ suggesting clients are motivated by the potential benefits or outcomes tied to their goals. ‘Generating Alternative Solutions’ is another significant strategy (20.03%), with words such as “brainstorming ideas” and “strategies tactics” showing the focus on exploring different options to address issues. The LIWC indicators ‘insight’ and ‘curiosity’ align with this focus on creativity and open-mindedness. Similarly, ‘Problem-Solving Positive Mindset’ bigrams, such as “positive attitude” and “healthy thinking”, show therapists’ encouragement to adopt a more constructive outlook on problems, fostering optimism. The prevalence of ‘emotion’ and ‘feeling’ reflects the emotional engagement involved in developing such positive mindset. ‘Outcome Prediction and Planning’ phrases like “pros cons” and “solve problem” indicate a focus on helping clients think through potential consequences. This is echoed by the prevalence of ‘risk’ and ‘focus future’ as dominant features, reinforcing the importance of careful consideration and future-oriented thinking.
On the ‘Communication’ side, the bigrams highlight how therapists build rapport and guide session flow. In ‘Therapeutic Engagement’, phrases like “really good” use affirming language to foster a supportive, collaborative atmosphere, helping clients feel understood. ‘Social Courtesies,’ with phrases like ‘thank much’ and ‘look forward,’ add warmth to interactions, maintaining a respectful and friendly tone. The top LIWC features, ‘emotion’ and ‘prosocial,’ reflect the positive social behaviors therapists use to foster mutual respect and connection. Meanwhile, in ‘Problem-Solving Test Review’, more formal words like “positive rational” and “impulsive avoidant” suggest an analytical approach, where therapists review the client’s PST test results in a structured, evaluative manner (See Appendix F for more information).
Progression of Strategies Across Therapy Visits. Figures 3 and 3 show a clear progression of strategies across three different visits of caregivers. In Figure 3, early sessions focus heavily on enhancing ‘Problem-Solving Mindset’ (Step One) and ‘Defining Problems and Goals’ (Step Two), indicating that the initial phase of therapy is centered on improving caregivers’ mindsets and helping them clearly define the challenges they face. As therapy progresses into the second and third visits, there is a shift towards more advanced strategies such as ‘Generating Alternative Solutions’ (Step Three) and ‘Outcome Prediction and Planning’ (Step Four), showing that the focus moves towards exploring possible solutions and planning for their implementation. By the third visit, ‘Trying Out Solution Plan’ (Step Five) becomes more prominent, suggesting that the final phase of therapy is geared towards applying and testing these solutions in real-life situations. Figure 3 shows the distribution of ‘Communication’ strategies across visits. Strategies such as ‘Therapeutic Engagement’, ‘Session Management’, and ‘Social Courtesies’, were applied consistently throughout the sessions. This indicates that these strategies are essential components of every visit, regardless of the stage of therapy. However, ‘PS Test Review’ was used more frequently during the first visit, reflecting the therapist’s focus on reviewing the client’s Problem-Solving Test results, which were completed before the therapy began. This early emphasis helps set the foundation for the sessions that follow. These intuitive results support the effectiveness of our classifier.
5 Discussion
Annotating PST-Related Texts. Our results show that GPT-4o significantly outperformed both GPT-4 and various open-weight models in annotating therapeutic strategies, achieving F1 score of 0.76 without the need for supplemental context. Despite Yi exhibiting the lowest entropy among the models tested, GPT-4o displayed higher consistency and reliability. This highlights critical trade-offs to consider between performance consistency and overall model reliability when deploying LLMs in sensitive fields such as mental health care. While GPT-4o demonstrates strong effectiveness, its success highlights the need for advancements in open-source models. Enhancing these models’ accuracy and reliability is essential to providing competitive and accessible alternatives in the healthcare domain. Such advancements could democratize access to high-quality mental health resources, thereby broadening the impact and reach of therapeutic interventions. Our models overall show better results compared to previous work Chiu et al. (2024), which reported that their best-performing classifier, GPT-4 with definitions and examples, achieved an F1 score of 57.7 in identifying therapist behavior across 13 different therapist codes.
Figure 2: Distribution of PS core Strategies in Different Visits
Figure 3: Distribution of Communication Strategies in Different Visits
Comparing the performance of the models in labeling ‘PS core’ vs. ‘Communication’ strategies, GPT-4o demonstrated notable efficacy in labeling structured, goal-oriented ‘PS core’ strategies such as ‘Defining Problems and Goals’ and ‘Generating Alternative Solutions,’ achieving high precision and recall. This performance highlights LLMs’ strength in recognizing structured, task-oriented language that is clear and objective.
On the other hand, GPT-4o was less robust when dealing with ‘Communication’ strategies like ‘Session Management’ and ‘Social Courtesies.’ These strategies are more dynamic and rely heavily on emotional and social cues. This disparity in performance highlights a key challenge for LLMs in therapeutic settings: while they excel at handling clear, problem-solving tasks, they struggle with the subtleties of interpersonal communication essential in therapy, lacking the ability to accurately interpret the nuances of human emotions and social interactions. These characteristics can often be implicit and vary significantly across different individuals and situations. Enhancing LLMs’ sensitivity to these subtle dynamics is important for improving their effectiveness within therapeutic environments. Research has indicated that LLMs can sometimes exhibit an over-realistic sense of empathy Sorin et al. (2023), stemming from a fundamental misunderstanding of underlying emotional contexts. To address this issue, future developments should concentrate on refining LLMs by incorporating strategies such as expert-in-the-loop systems Afzal et al. (2024); Sun et al. (2023). These systems should allow for real-time feedback and guidance from human experts, which could enhance the models’ ability to recognize and interpret complex emotional and social cues.
Model Fine-Tuning and Generalizability. The results of our fine-tuned transformer-based models showed that MentalBERT performed best in annotating problem-solving core strategies with an F1 score of 0.78, while DeBERTa excelled in communication strategies with a 0.73 F1 score. This indicates that different models can specialize in distinct aspects of therapeutic dialogue. Interestingly, our transformer-based model achieved an overall F1 score of 0.68, comparable to our second-best LLM.This result is particularly noteworthy as it suggests that both MentalBERT and DeBERTa, with their distinctive capabilities, can serve as reliable and accessible annotators for a range of therapeutic strategies, presenting a promising avenue for scalable and efficient analysis of therapy content. These models can also be more appropriate in the context of therapy where private and sensitive information is being analyzed Shah et al. (2022). This capability is important for maintaining confidentiality and trust in digital therapeutic environments, which is fundamental for effective therapy.
Real-World Applications of LLM-Annotated Data. The strong performance of our models, especially GPT-4o, in annotating therapeutic strategies, even without conversational context, demonstrates the potential for using LLMs to automate the analysis of PST transcripts at scale. By integrating LLMs into real-time therapy support or documentation systems, therapists could receive automated insights that help them refine their approaches, track client progress, and identify appropriate strategies more efficiently (As shown in Figures 3 and 3). These systems could also tailor interventions to individual client needs and offer real-time suggestions to therapists, further optimizing therapeutic outcomes. However, this raises important questions about the risks of over-reliance on AI-generated feedback. Our findings indicate that while LLMs are adept at processing and analyzing data, they sometimes fail to accurately identify specific therapeutic strategies and lack the capability to convey genuine empathy and emotion. While LLMs can provide valuable assistance, ensuring that human judgment remains central to the therapeutic process is crucial, as LLMs may struggle with the nuanced emotional and interpersonal dynamics that are critical in therapy. Balancing AI support with therapist oversight will be key to maximizing the benefits of these technologies while minimizing potential risks.
6 Conclusion
This study highlights the potential of LLMs like GPT-4o to enhance PST through effective annotation of therapeutic dialogues. GPT-4o achieved a notable F1 score of 0.76 in the ‘No Context’ condition, with a low entropy value of 0.035, demonstrating its efficiency and reliability in categorizing therapeutic strategies. Additionally, we fine-tuned transformer-based models on a sample of 5,000 therapist utterances annotated by GPT-4o. Notably, DeBERTa matched the performance of our second-best LLM, GPT-4, showcasing the effectiveness of integrating AI with human expertise. We also expanded our codebook to include a new ‘Communication’ dimension, enriching our analysis of therapist-client interactions. Our findings showed that all models performed better at identifying ‘PS core’ strategies, indicating a strong capability in recognizing structured therapeutic elements. This research sets the stage for future innovations that enhance the integration of technology in mental health care, making therapeutic interventions more accessible and impactful.
7 Limitations
Our study has several limitations, including the dependence on ‘transcribed’ therapy sessions, which may not fully capture the dynamic aspects of live interactions such as tone and non-verbal cues. The use of LLMs like GPT-4o and GPT-4 introduces potential biases from their internet-based training data, impacting the interpretation of therapeutic strategies. Furthermore, our analysis is based solely on English-language data, which limits the model’s applicability in non-English-speaking contexts. These limitations highlight the technology’s current stage and its focus on textual analysis, excluding the broader emotional and psychological facets of therapy sessions.
In addition, we only focus on PST-related framework and analysis which may not be generalizable to other therapeutic methods like cognitive-behavioral therapy (CBT) or interpersonal therapy (IPT). These methods may involve different communication styles and dynamics that the LLMs we used may not accurately interpret. Future work will analyze the adaptability of our model for broader therapy analysis. Additionally, cultural and contextual differences in therapy could further limit the model’s applicability across diverse mental health settings. Future research could address this by testing the models in a broader range of therapeutic approaches and contexts.
Finally, we did not include any analysis of LLM-generated therapy data, which could offer a valuable comparison between real-world therapeutic interactions and simulated data produced by LLMs. Our focus is exclusively on the characteristics of authentic, real-world data from therapy sessions. Future research could explore this comparison, examining how well LLM-generated interactions align with or diverge from real-world therapeutic PST dialogues, potentially offering insights into improving the practicality and effectiveness of AI-driven interventions in mental health care.
8 Ethical Considerations
A few of the coauthors provided the deidentified transcriptions of PST sessions, which had been collected as part of a large trial assessing the comparisons across different modalities of PST delivery. We used these transcriptions to conduct secondary analysis and experiment on automated annotation, as we have presented in this paper. The institutional review boards have determined this study as not human subjects research, and a data-sharing agreement has been signed.
The application of LLMs in healthcare, particularly in sensitive areas like cognitive behavioral therapy, necessitates careful consideration of their implications. Ensuring responsible development and use of AI technologies in mental health care is crucial, requiring vigilance for biases, stakeholder engagement, and strong safeguards to maintain individuals’ confidentiality and the integrity of the therapeutic process. This paper does not endorse the use of LLMs in therapeutic settings, nor does it claim that they are ready for such applications. Rather, our goal is to systematically evaluate the behavior of current LLMs when applied in therapeutic contexts. The use of AI in mental health care raises ethical concerns about automating therapy. AI tools, like LLMs, should assist rather than replace therapists, ensuring they do not undermine the personal, empathetic nature of therapy, which is crucial for effective treatment. While AI can provide useful insights, therapists must remain the primary decision-makers, as AI cannot fully grasp the complexities of human behavior and emotions as shown in our results and other related work. Accountability should always rest with human professionals to preserve the integrity and effectiveness of mental health care.
9 Acknowledgment
We are deeply grateful to Dr. Bum Chul Kwon, Max Song, and DongWhan Jun for their invaluable insights and assistance throughout this project. This paper is based upon work supported by the National Science Foundation under Grant #2144880.
Appendix A Appendix
Appendix B Therapists Strategies
Table 6 presents the final list of therapist strategies for PST, with definitions and examples for each category.
Category Strategy Definition Example
Problem-solving Positive Mindset Strategies building a positive approach towards challenges, emphasizing optimism and emotional awareness. “uh, they suggest trying to think of problems as challenges rather than threats.”
Defining Problems and Goals Setting achievable goals and identifying factual problems while recognizing obstacles. “Are there any obstacles to those goals? Can we list anything that is getting in the way?”
PS Core Generating Alternative Solutions Brainstorming solutions, differentiating between strategies, and encouraging open idea exploration. “So let’s brainstorm some specific strategies and tactics for how you can make that happen.”
Evaluating Outcomes and Developing a Plan Narrowing down solutions, assessing impacts, and creating an action plan. “Working through this worksheet…short-term personal consequences of self-care are pretty positive, right?”
Trying Out Solution Plan Implementing and assessing the solution strategy, troubleshooting, and acknowledging successes. “Then just monitoring your outcomes. Has the problem improved in some ways?”
Social Courtesies* Polite, surface-level interactions like greetings, small talk, and farewells to create a friendly atmosphere. “thank you again for your time today, and have a really good week.”
Communication Session Management* The therapist provides guidance on session logistics, manages time, keeps discussions on track, and helps plan future sessions. “So you got Skype up and running, and that is so fun!”
Therapeutic Engagement* The therapist shows attention, validates emotions, offers support, reflects feelings, praises strengths, and provides comfort or personal insight. “I’m glad it’s starting to show some small improvements.”
Problem-Solving Test Review* The therapist explains the PS Test and reviews the client’s results across five dimensions: positive/negative problem orientation, rational, impulsive, and avoidant problem-solving styles. “So positive and rational were your highest scores. Your positive was, uh, really, is almost as high as it could possibly get.”
Table 6: Overview of Therapist Strategies for PST, *denotes newly created codes
Appendix C Model Setting
LLM parameters:
temperature = 0;
max-tokens = 500;
openai-api-version=‘gpt-4o-2024-05-13’, ‘gpt-4-0125-preview’, ‘meta-llama/Llama-3.1-70B-Instruct’, ‘01-ai/Yi-1.5-34B-Chat’
Transformer-based Models Parameters:
Adam optimizer, learning-rate=2e-5, DeBERTa = microsoft/deberta-v3-base, MentalBERT = mental/mental-bert-base-uncased, FlanT5 = google/flan-t5-base
Appendix D Prompts
In this section, we present the specific prompt used to obtain annotations from LLMs (the few-shot examples are not included here).
D.1 Without Context
Instruction.
You are a highly helpful assistant tasked with annotating the therapist’s behavior. Your role is to categorize the therapist’s behavior using a list of strategies provided below.
Your Task: Select one strategy from the ‘PS core’ category and one strategy from the ‘Communication’ category that best aligns with the therapist’s utterance. If none of the strategies are applicable, return ’None’ for that category.
Definition.
Definitions of Therapist’s Strategies for Problem-Solving Therapy:
PS core Strategies: 1- Problem-solving Positive Mindset: Techniques that foster a positive outlook, such as promoting optimism, reframing negative thoughts, visualizing success, practicing gratitude, mindfulness, positive self-talk, and emotional awareness. 2- Defining Problems and Goals: Identifying clear, achievable goals, recognizing obstacles, and determining which specific issues to address in therapy. 3- Generating Alternative Solutions: Brainstorming a variety of solutions without judgment, distinguishing between strategies and tactics, and encouraging open exploration of ideas. 4- Outcome Prediction and Planning: Narrowing down solutions, evaluating their impact (personal, social, temporal), and creating a detailed action plan for implementation. 5- Trying Out Solution Plan: Implementing the solution, assessing its effectiveness, troubleshooting issues, seeking support, and acknowledging successes in the process.
Communication Strategies: 6- Social Courtesies: Engaging in polite, yet somewhat superficial social interactions like greetings, small talk (e.g., asking about the weather or weekend plans), thanking the client for their time, and saying goodbye. 7- Session Management: Managing session logistics, such as informing the client about session length, guiding the direction of the conversation, refocusing when off-topic, or scheduling future sessions. 8- Therapeutic Engagement: Actively listening, validating the client’s feelings, providing emotional support, recalling previous information shared by the client, reflecting emotions, normalizing experiences, commenting on client strengths, offering personal disclosure, and offering comfort. 9- Problem-solving Review Test: Discussing the client’s Problem-Solving Test (PST) results, explaining the test, and reviewing five problem-solving dimensions: positive problem orientation, negative problem orientation, rational style, impulsive style, and avoidant style.
D.2 With Context
Instruction.
You are a highly helpful assistant tasked with annotating the therapist’s behavior. Your role is to categorize the therapist’s behavior using a list of strategies provided below.
Your Task: Select one strategy from the ‘PS core’ category and one strategy from the ‘Communication’ category that best aligns with the therapist’s most recent utterance. You are also provided with the last two utterances: the therapist’s previous utterance and the client’s previous utterance. You should use only these two previous utterances as context, but for annotation, you should focus only on the most recent therapist’s utterance. If none of the strategies are applicable, return ’None’ for that category.
Definition.
The definition is the same as when context is not provided.
Appendix E Detailed Results of Classification
E.1 LLM Classification Results
The detailed classification results for each strategy, based on the F1-scores of different LLMs, are presented in Table 7.
Strategies GPT-4 GPT-4o Llama Yi
PS Core Problem-solving Positive Mindset 0.83 0.85 0.76 0.75
Defining Problems and Goals 0.73 0.79 0.72 0.63
Generating Alternative Solutions 0.74 0.79 0.62 0.48
Outcome Prediction and Planning 0.74 0.84 0.67 0.50
Trying Out Solution Plan 0.89 0.80 0.75 0.40
Communication Social Courtesies 0.75 0.73 0.58 0.67
Session Management 0.72 0.68 0.61 0.61
Therapeutic Engagement 0.57 0.80 0.69 0.64
Problem-solving Test Review 0.89 0.92 0.79 0.70
None 0.61 0.62 0.28 0.27
weighted avg 0.68 0.76 0.61 0.56
Table 7: F1-scores of different LLMs for each strategy (without context)
E.2 Fine-tuned Models Classification Results
The detailed classification results for each strategy, based on the F1-scores of different fine-tined transformer-based models, are available in Table 8.
Strategies DeBERTa MentalBERT FlanT5
PS Core Problem-solving Positive Mindset 0.81 0.78 0.76
Defining Problems and Goals 0.80 0.79 0.79
Generating Alternative Solutions 0.73 0.70 0.75
Outcome Prediction and Planning 0.73 0.77 0.70
Trying Out Solution Plan 0.80 0.89 0.40
Communication Social Courtesies 0.73 0.67 0.68
Session Management 0.66 0.63 0.62
Therapeutic Engagement 0.70 0.68 0.69
Problem-solving Test Review 0.83 0.79 0.80
None 0.41 0.53 0.50
weighted avg 0.68 0.68 0.67
Table 8: F1-scores of different fine-tuned models for each strategy
Appendix F LIWC Results
Figures 4 and 5 present the co-occurrence of each strategy with the selected LIWC features.
Figure 4: Co-Occurrence of PS core Strategies with Selected LIWC Features
Figure 5: Co-Occurrence of Communication Strategies with Selected LIWC Features